# Improving large language models with concept-aware fine-tuning

### v1

@@ 3.2 Downstream Tasks
We present the results on five downstream tasks of different domains to demonstrate the superior performance and broad applicability of concept-aware fine-tuning. Example questions from every evaluation dataset can be found in Figure 2b. The general experimental setup is as follows: using task-specific training datasets, we fine-tune four separate models using full CAFT, LoRA CAFT, and their next-token versions. The latter two serve as baseline models for comparison. They are trained for up to 5 epochs with early stoppage and for sequence lengths ranging from 512 to 2048, depending on the specific task. All results shown are the average of 5 independent evaluation runs. From Task 3 onwards, the auxiliary heads are pretrained on the training set for 1 epoch before proceeding to the actual fine-tuning.
@@ 3.2.1 Downstream Task 1: Coding
HumanEval Base 38.9
LoRA Fine-tuning Next-token 40.9
CAFT 45.1
Full Fine-tuning Next-token 40.5
CAFT 49.3
@@ 3.2.2 Downstream Task 2: Mathematics
MATH-500 Base 19.1
LoRA Fine-tuning Next-token 22.9
CAFT 24.6
Full Fine-tuning Next-token 23.7
CAFT 25.2
@@ 3.2.3 Downstream Task 3: Text Generation
Base 29.17 6.64 15.46 27.49
LoRA Fine-tuning Next-token 42.31 20.56 29.86 40.44
CAFT 44.16 22.3 31.62 42.37
Full Fine-tuning Next-token 44.57 22.94 32.17 42.75
CAFT 45.93 24.44 33.76 44.04
@@ 3.2.4 Downstream Task 4: Molecular Generation
For this case study, the train and test sets are drawn from the challenging L+M-24 dataset (Edwards et al., 2024). The input is a natural language description of the desired molecular properties, and the output is the associated molecule in the form of a SMILES sequence. The following metrics are used: (i) exact match (EM), representing the proportion of generated SMILES sequence that exactly matches the gold truth, (ii) Levenshtein distance (Miller et al., 2009), measuring the statistical distance between the model generation and ground truth, (iii) Morgan Fingerprint Tanimoto Similarity (FTS) (Rogers, Hahn, 2010), measuring the structural similarity between SMILES strings, and finally (iv) validity, representing the proportion of generations that conform to SMILES grammar and chemical rules.
@@
Task Method EM (%) $\uparrow$ Levenshtein $\downarrow$ Morgan FTS $\uparrow$ Validity (%) $\uparrow$
L+M-24 Base 0.00 139.19 4.71 68.70
LoRA Fine-tuning Next-token 0.00 64.14 62.42 91.84
CAFT 0.00 58.69 63.20 93.46
Full Fine-tuning Next-token 0.14 47.59 65.34 92.38
CAFT 0.54 41.23 67.79 97.14
@@ 3.2.5 Downstream Task 5: De Novo Protein Design
Base 12.06 -71.59 22.28 6.80
Full Fine-tuning Next-token 20.32 -16.01 52.6 33.07
CAFT 22.14 3.18 54.3 35.12
@@ 4 Related Works
Downstream Fine-tuning. Importantly, foundational models are often further trained by industry practitioners and researchers to perform domain-specific tasks, such as reasoning (Guo et al., 2025), math (Liu et al., 2024b), and even molecular generation (Fang et al., 2023). However, because fine-tuning all model parameters may be too computationally expensive, parameter-efficient fine-tuning methods such as LoRA (Hu et al., 2022) and QLoRA (Dettmers et al., 2023) were introduced.
@@ 5.1 The Advantages of Concept-Aware Fine-Tuning (CAFT)
Concept-aware fine-tuning (CAFT) addresses LLMs’ fragmented conceptual understanding, caused by the artificial nature of tokenization’s text parsing. Specifically, CAFT is the first method to introduce multi-token prediction to model fine-tuning. This was previously thought to be unfeasible due to the brevity of the post-training phase. Empirically, CAFT leads to substantial performance gains by improving concept formation.
@@
5.2 Broader Implications: Rethinking the next-token objective
LLMs’ ability to plan beyond the next token has long been controversial. Learning across tokens is crucial for coherent text generation and accomplishing a broad spectrum of tasks. The broad success of next-token LLMs, starting from GPT-3 (Brown et al., 2020), strongly suggests that existing models can learn coherent concepts across tokens to some extent. Other empirical works confirm this: for one, Lindsey et al. (2025) find that LLMs plan beyond the next word when writing poems.
@@ 5.2 Broader Implications: Rethinking the next-token objective
Thus far, multi-token training has been held back by being restricted to the pretraining phase and by its prohibitive computation costs. Given that CAFT directly addresses these factors, we hope the introduction of CAFT will drive the adoption of the multi-token setting during post-training, potentially becoming the de facto training objective for model fine-tuning.
@@ Appendix B Understanding the CAFT hyperparameters
The introduction of the CAFT hyperparameters $\alpha$, $\beta$, and $\gamma$ is key to CAFT’s effectiveness. Specifically, they ensure that model’s leverage the additional information provided by the auxiliary heads, while ensuring they focus on optimizing $\mathcal{L}_{1}$. Figure 5 visualizes how they scale auxiliary losses over iterations, assuming $\alpha=0.8$, $\beta=0.01$, and $\gamma$ is defined by the reflected sine schedule, which is formally:
@@ C.2 Full and LoRA Concept-Aware Fine-Tuning (CAFT)
For the CAFT-specific hyperparameters, the objective is to incentivize models to leverage the auxiliary losses and ultimately optimize for $\mathcal{L}_{1}$. We conducted an extensive hyperparameter search and found the following settings to be robust across all tasks. $\gamma$ is defined using a reflected sine schedule. Empirically, we find that this performs better than a constant or sine schedule. Other hyperparameters are shown in Table 7. Additionally, when doing a hyperparameter search, we find that $\alpha=0.8$ performs better than 0.7 and 0.9, while $\beta=0.01$ performs better than 0.05 and 0.10. In practice, these values do not need to be further tuned.
@@ C.3 Evaluation
All base, next-token fine-tuned, and CAFT models are provided a zero-shot prompt that gives explicit instructions on the desired output format. Models are asked to generate the relevant code, text, or sequence only, except for MATH-500, where models are asked to “think step-by-step". To control for noise, 5 independent runs are done for every dataset; their average metrics are reported. Given that early stoppage is used, models are also evaluated on earlier checkpoints, and the top-performing one is reported.

### v2

@@ 3.2 Downstream Tasks
We present the results on five downstream tasks of different domains to demonstrate the superior performance and broad applicability of concept-aware fine-tuning. Example questions from every evaluation dataset can be found in Figure 2b. The general experimental setup is as follows: using task-specific training datasets, we fine-tune four separate models using full CAFT, LoRA CAFT, and their next-token versions. The latter two serve as baseline models for comparison. They are trained for up to 5 epochs with early stoppage and for sequence lengths ranging from 512 to 2048, depending on the specific task. All results shown are the average of 5 independent evaluation runs and their variance via 95% confidence intervals (CIs). From Task 3 onwards, the auxiliary heads are pretrained on the training set for 1 epoch before proceeding to the actual fine-tuning.
@@ 3.2.1 Downstream Task 1: Coding
HumanEval Base 38.9  $\pm$1.501
LoRA Fine-tuning Next-token 40.9  $\pm$0.927
CAFT 45.1$\pm$1.930
Full Fine-tuning Next-token 40.5  $\pm$2.309
CAFT 49.3$\pm$2.590
@@ 3.2.2 Downstream Task 2: Mathematics
MATH-500 Base 19.1  $\pm$1.367
LoRA Fine-tuning Next-token 22.9  $\pm$1.033
CAFT 24.6$\pm$1.666
Full Fine-tuning Next-token 23.7  $\pm$0.572
CAFT 25.2$\pm$0.732
@@ 3.2.3 Downstream Task 3: Text Generation
Base [0pt] [0pt] 29.17 (0.188) 6.64 (0.015) 15.46 (0.048) 27.49 (0.160)
LoRA Fine-tuning Next-token [1pt] [0pt] 42.31 (0.173) 20.56 (0.136) 29.86 (0.150) 40.44 (0.133)
CAFT [1pt] [0pt] 44.16 (0.277) 22.30 (0.212) 31.62 (0.176) 42.37 (0.288)
Full Fine-tuning Next-token [1pt] [0pt] 44.57 (0.219) 22.94 (0.231) 32.17 (0.272) 42.75 (0.218)
CAFT [1pt] [0pt] 45.93 (0.315) 24.44 (0.281) 33.76 (0.306) 44.04 (0.327)
@@ 3.2.4 Downstream Task 4: Molecular Generation
For this case study, the train and test sets are drawn from the challenging L+M-24 dataset (Edwards et al., 2024). The input is a natural language description of the desired molecular properties, and the output is the associated molecule in the form of a SMILES sequence. The following metrics are used: (i) exact match (EM), representing the proportion of generated SMILES sequence that exactly matches the gold truth, (ii) Levenshtein distance (Miller et al., 2009), measuring the statistical distance between the model generation and ground truth, (iii) RDK Fingerprint Tanimoto Similarity (FTS) (Schneider et al., 2015), measuring the structural similarity between SMILES strings, and finally (iv) validity, representing the proportion of generations that conform to SMILES grammar and chemical rules.
@@
Method EM (%) $\uparrow$ Levenshtein $\downarrow$ RDK FTS $\uparrow$ Validity (%) $\uparrow$
Base [0pt] [0pt] 0.00 (0.000) 139.19 (3.621) 9.67 (0.383) 68.70 (1.563)
LoRA Fine-tuning Next-token [1pt] [0pt] 0.00 (0.000) 64.14 (3.284) 62.42 (0.178) 91.84 (1.019)
CAFT [1pt] [0pt] 0.00 (0.000) 58.69 (3.661) 63.20 (0.362) 93.46 (1.158)
Full Fine-tuning Next-token [1pt] [0pt] 0.14 (0.068) 47.59 (2.565) 65.34 (0.126) 92.38 (0.778)
CAFT [1pt] [0pt] 0.54 (0.068) 41.23 (0.801) 67.79 (0.172) 97.14 (0.655)
@@ 3.2.5 Downstream Task 5: De Novo Protein Design
Base [0pt] 12.06 (0.219) -71.59 (3.680) 22.28 X 6.80 X
Full Fine-tuning Next-token [1pt] [0pt] 20.32 (0.180) -16.01 (0.208) 52.60 X 33.07 X
CAFT [1pt] [0pt] 22.14 (0.180) 3.18 (1.369) 54.30 X 35.12 X
@@ 4 Related Works
Downstream Fine-tuning. Importantly, foundational models are often further trained by industry practitioners and researchers to perform domain-specific tasks, such as math (Liu et al., 2024b), reasoning (Chen et al., 2025), and even molecular generation (Fang et al., 2023). However, because fine-tuning all model parameters may be too computationally expensive, parameter-efficient fine-tuning methods such as LoRA (Hu et al., 2022) and QLoRA (Dettmers et al., 2023) were introduced.
@@
Concept-based Learning. Several works have sought to incorporate concepts into neural networks. Concept bottleneck models are trained to predict human-labeled concepts, which are then used for final prediction; these are typically used for computer vision tasks (Koh et al., 2020; Kumar et al., 2009). The Large Concept Model (Barrault et al., 2024), defining sentences as concepts, performs next-sentence prediction in an embedding space. Finally, recent works on training LLMs to reason in continuous latent spaces (Hao et al., 2024; Zhang et al., 2025) also draw inspiration from conceptual thinking.
@@ 5.1 The Advantages of Concept-Aware Fine-Tuning (CAFT)
Concept-aware fine-tuning (CAFT) addresses LLMs’ fragmented conceptual understanding, caused by the artificial nature of tokenization’s text parsing. Specifically, CAFT is the first method to introduce multi-token prediction to model fine-tuning. This was previously thought to be unfeasible due to the dramatic distribution shift and brevity of the post-training phase. CAFT resolves these problems by introducing two main novel techniques: (i) pretraining the auxiliary heads on a general instruction tuning dataset and (ii) dynamically scaling the auxiliary losses. Empirically, CAFT leads to substantial performance gains by improving concept formation.
@@
5.2 Broader Implications
Rethinking the next-token objective. LLMs’ ability to plan beyond the next token has long been controversial. Learning across tokens is crucial for coherent text generation and accomplishing a broad spectrum of tasks. The broad success of next-token LLMs, starting from GPT-3 (Brown et al., 2020), strongly suggests that existing models can learn coherent concepts across tokens to some extent. Other empirical works confirm this: for one, Lindsey et al. (2025) find that LLMs plan beyond the next word when writing poems.
@@ 5.2 Broader Implications
Thus far, multi-token training has been held back by its restriction to the pretraining phase and its prohibitive computation costs. Given that CAFT directly addresses these factors, we hope the introduction of CAFT will drive the adoption of the multi-token setting during post-training, potentially becoming the de facto training objective for model fine-tuning.
Tokenization. As LLMs continue to evolve, token vocabulary sizes have only increased. This is because larger vocabularies improve the representation of unique, domain-specific text fragments and reduce the splitting of words into subwords. For example, from Llama 2 to 3 (Grattafiori et al., 2024), vocabulary sizes have increased by four times, from 32K to 128K. However, this exponentially increases model sizes and training costs; a token size of 128K causes the unembedding layer to balloon to over 500M parameters. Other tokenization methods have, in parallel, sought to capture coherent word phrases (Lai et al., 2021; Liu et al., 2025).
We posit that CAFT reduces the reliance on specific tokenization strategies. Through multi-token prediction, CAFT allows models to capture coherent phrases even when each independent token is an arbitrary, incoherent subword. Optimal vocabulary sizes may also be lower than prescribed by existing scaling laws (Tao et al., 2024). Empirical studies of these ideas can be pursued in future works.
@@ Appendix B Understanding the CAFT hyperparameters
The introduction of the CAFT hyperparameters $\alpha$, $\beta$, and $\gamma$ is key to CAFT’s effectiveness. Specifically, they ensure that model’s leverage the additional information provided by the auxiliary heads, while ensuring they focus on optimizing $\mathcal{L}_{1}$. Figure 5 visualizes the loss scaling of each auxiliary head loss over 1000 iterations, assuming $\alpha=0.8$, $\beta=0.01$, and $\gamma$ is defined by the reflected sine (or RSine) schedule, which is formally:
@@ C.2 Full and LoRA Concept-Aware Fine-Tuning (CAFT)
CAFT $\gamma$ RSine - RSine
For the CAFT-specific hyperparameters, the objective is to incentivize models to leverage the auxiliary losses and ultimately optimize for $\mathcal{L}_{1}$. We conducted an extensive hyperparameter search and found the following settings to be robust across all tasks. $\gamma$ is defined using a reflected sine (or RSine) schedule. Empirically, we find that this performs better than a constant or sine schedule. Other hyperparameters are shown in Table 7. Additionally, when doing a hyperparameter search, we find that $\alpha=0.8$ performs better than 0.7 and 0.9, while $\beta=0.01$ performs better than 0.05 and 0.10. In practice, these values do not need to be further tuned.
@@ C.3 Evaluation
All base, next-token fine-tuned, and CAFT models are provided a zero-shot prompt that gives explicit instructions on the desired output format. Models are asked to generate the relevant code, text, or sequence only, except for MATH-500, where models are asked to “think step-by-step". Given that early stoppage is used, models are also evaluated on earlier checkpoints, and the top-performing one is reported.
To control for noise, 5 independent runs are done for every dataset; their average metrics are reported. Additionally, following Grattafiori et al. (2024), we report each score’s variance via 95% confidence intervals (CIs). The confidence intervals of pLDDT and TM-score in Table 5 are not reported because the highest scoring output from ColabFold is used instead of the average, as is standard practice in the existing literature.
@@
C.4 Experimental Setup of Conceptual Understanding Empirical Study (Section 3.3)
To extract multi-token concepts from the model generations, we first conducted a qualitative analysis to identify common ways in which concepts are expressed in the ground truth Python solutions in HumanEval. We found the following types of text fragments to be reasonable proxies: (i) text contained within brackets, e.g. (1 + 1), (ii) text contained within quotations, e.g. "__main__", and (iii) objects separated by periods, e.g. json.loads. These fragments are extracted via a Python parser; the number of fragments corresponds to the number of concepts in a question. Questions are then split into two bins, separated by the mean concepts per question of 5.
As for molecular generation, we define functional groups as chemical concepts within a given molecular structure. Benzene, amide, and carboxylic acid groups are chosen for their relatively long sequence lengths and sufficient frequency in the dataset. The matched (%) metric is calculated based on whether the model generation contains the same number of a given functional group as the ground truth, given that the ground truth contains at least 1 group. To account for false positives (model generates functional groups for questions that do not contain them), the F1 score is also reported.